# SeaMo: A Multi-Seasonal and Multimodal Remote Sensing Foundation Model

### v1

@@ Abstract
Remote Sensing (RS) data contains a wealth of multi-dimensional information crucial for Earth observation. Owing to its vast volume, diverse sources, and temporal properties, RS data is highly suitable for the development of large Visual Foundation Models (VFMs). VFMs act as robust feature extractors, learning from extensive RS data, and are subsequently fine-tuned for deployment in various geoscientific tasks. However, current VFMs in the RS domain are predominantly pretrained and tailored exclusively for specific characteristics of RS imagery, neglecting the potential of utilizing the multi-dimensional properties of RS data. Therefore, in this work, we propose SeaMo, a pioneering visual foundation model that integrates multi-seasonal and multimodal information in the RS field. SeaMo is designed to harness multiple properties of RS data. Within the masked image modeling framework, we employ non-aligned cropping techniques to extract spatial properties, use multi-source inputs for multimodal integration, and incorporate temporal-multimodal fusion blocks for effective assimilation of multi-seasonal data. SeaMo explicitly models the multi-dimensional properties of RS data, making the model more comprehensive, robust, and versatile. We applied SeaMo to several downstream geoscience tasks, which demonstrated exceptional performance. Extensive ablation studies were conducted to validate the model’s superiority.
@@ 1 Introduction
Remote sensing is a technology that enables the measurement of soil or crop characteristics using platforms such as UAVs, airplanes, or satellites [1]. The data acquired through remote sensing technology are utilized to study social phenomena, land use dynamics, and changes in landscapes for scientific analysis and modeling purposes. Currently, hundreds of remote sensing satellites continuously monitor the Earth’s surface, generating large-scale time series datasets. These datasets, characterized by their heterogeneity and incorporation of multiple sources, including spectral data, radar data, and meteorological data, provide a rich repository of physical geographic information, embodying the typical traits of big data [2]. The research on image data in remote sensing currently faces several challenges and focal points:
@@
(1) Integration of data from diverse sources for varied geoscientific applications: Different methods of capturing how electromagnetic radiation interacts with Earth surface materials provide unique datasets [3]. Optical data, by measuring radiation across many wavelengths, offers detailed spectral information that characterizes the material composition of objects. Synthetic Aperture Radar (SAR) data, by transmitting and receiving longer wavelength electromagnetic pulses, can assess the geometry, roughness, and electrical properties of objects. Given the significant heterogeneity between these data sources, simple transformations and combinations of data dimensions are insufficient to fully exploit this information [4]. Thus, a key area of focus is how to extract valuable geoscientific information from these diverse datasets that can meet human needs.
@@
(2) Effective modeling and utilization of spatio-temporal data: The spatio-temporal structure of satellite imagery involves the availability of multiple temporally-spaced images for the same location. This rich spatio-temporal characteristic of remote sensing data enables the analysis of changes in physical geography, and land use, and offers insights into geological structures or socio-economic conditions [5]. However, the low information density of spatio-temporal data, due to low spatial resolution and long intervals between observations, makes it challenging to capture temporal dynamics with simple techniques. Therefore, how to effectively mine geoscientific information from spatio-temporal data with low information density remains a research hotspot.
@@
Visual Foundation Models (VFMs) refer to large-scale neural networks that are pretrained on vast datasets, typically using self-supervised learning methods. VFMs act as robust and powerful feature extractors and can be fine-tuned with minimal effort for various downstream tasks [6]. Given the inherent big data characteristics and the complex information contained within remote sensing data, the vast data landscape of remote sensing offers significant opportunities for the implementation of VFMs in earth observation. Currently, several VFMs have been introduced in the field of remote sensing, demonstrating performance that surpasses traditional deep learning models across a broad range of geoscientific tasks. This advancement has led to significant breakthroughs in the mining and extraction of remote sensing information. However, these models still face several challenges and deficiencies:
@@
(1) Lack of geoscientific attributes: Some VFMs in the remote sensing domain are derived from computer vision technologies and often involve simplistic transformations of remote sensing data to adapt it to algorithmic requirements. This approach generally leads to insufficient exploration of the geoscientific properties inherent in remote sensing data, resulting in a gap in feature extraction capabilities. For instance, DINO-MC [7] employs basic data augmentation and uses Contrastive Learning (CL) for pretraining. Similarly, RSVA [8] straightforwardly adapts the Masked Autoencoders from the computer vision domain for pretraining in remote sensing.
@@
(2) Single-dimension data modeling: While most VFMs acknowledge the unique characteristics of remote sensing data, they typically focus on a single dimension, such as spatial or temporal attributes [9]. For instance, SeCo [10] and CaCo [11] exploit the temporal dynamics of RS images, employing Contrastive Learning (CL) to capture time-invariant features at specific geo-locations. RingMo [12] introduces Masked Image Modeling (MIM) to the RS domain, targeting small objects with a sparse masking strategy. SatMAE [13] employs a group combination masking strategy within the Masked Autoencoders (MAE) framework [14], enriching its handling of multi-spectral or temporal data. Similarly, SpectralGPT [15] innovatively addresses the spectral sequence attributes of images, employing a 3D mask strategy focused on spectral sequence reconstruction. ScaleMAE [16] and SatMAE++ [17] adjust to the varying sizes and resolutions of RS images, designing multiple image-size reconstruction targets. CROMA [18] pretrains aligned radar and optical images using both CL and MIM methods. The objective is to leverage the multimodal attributes of remote sensing to enhance the model’s feature extraction capabilities.
@@
(3) Non-explicit modeling multi-dimensional data attributes: Some VFMs consider multi-dimensional attributes, yet their approach to modeling and pretraining is more akin to stacking and assembling rather than explicitly modeling from a remote sensing and geoscientific perspective [19, 20]. This often leads to a superficial integration of multidimensional data, which does not fully exploit the synergistic potential of combining spatial, temporal, and spectral characteristics inherent in remote sensing data. For instance, Skysense [19] integrates information from various remote sensing (RS) data modalities, temporal information, and geographic location within the framework of contrastive learning. It employs contrastive learning for multimodal spatial feature extraction and utilizes positional encodings to embed temporal and geographic information, surpassing previous models. However, the integration of its multi-dimensional attributes is not fully realized. Specifically, the model optimizes the data learning from multiple separate aspects rather than in a holistic manner. The embedding of positional encodings serves merely as an insertion of prior knowledge rather than facilitating active knowledge learning and discovery.
@@
From the analysis of existing research questions and deficiencies in remote sensing visual foundation models, we believe it is necessary to establish a visual foundation model that not only effectively utilizes multi-source satellite data but also explicitly models the spatial and temporal properties of the data. Furthermore, this model should be optimized using a unified objective during training. Therefore, we introduce SeaMo, a pioneering foundation model in the RS field that effectively integrates spatial, multimodal, and multi-seasonal information. To enhance the spatial information density, we utilize unaligned crop techniques as data augmentation methods. To leverage the complementary strengths of multimodal satellite data during model pretraining, we employ a unified encoder that allows for multisource data, which is then jointly input into the model. The integration of information is facilitated through self-attention mechanisms. To enhance time-invariant representations, we have designed a temporal fusion block with cross-attention mechanisms that effectively integrates multiple spatial-temporal data streams. Finally, the model is optimized using a unified masked modeling approach. Our contributions are outlined as follows:
@@
    We propose a unified Masked Image Modeling (MIM) framework that learns representations from multimodal and multi-seasonal data. For multimodal data, we utilize a concat-style strategy, while for multi-seasonal data, we implement a siamese network approach.
@@
    We have developed a Temporal-Multimodal (TM) fusion block, a cascade-style component that effectively integrates different seasonal and multimodal information during pretraining.
@@
    Our pretraining methodology adopts a multi-stage strategy that incrementally progresses from learning unimodal representations to multimodal and then to seasonal-multimodal representations. This step-by-step approach ensures a comprehensive understanding of the multidimensional features of RS data.
@@
    The proposed foundation model has been successfully transferred to various downstream geoscientific applications across different modalities, demonstrating its robust performance. Additionally, detailed ablation studies have been conducted to further substantiate the superior capabilities of the model.
@@ 2.1 Self-supervised visual representation learning
Self-supervised learning is a technique for learning generalizable visual representations through various pretext tasks without the need for annotations during pretraining [21]. Initially, this approach utilized straightforward tasks such as image colorization, image rotation prediction, or solving jigsaw puzzles [22, 23, 24]. In recent years, Contrastive Learning (CL) and Masked Image Modeling (MIM) have become predominant self-supervised techniques for pretraining large foundation models. CL is designed to capture feature invariance across a batch of images by differentiating between identical and modified images using data augmentation techniques, with implementations like SimCLR [25] and MoCo [26]. Conversely, MIM focuses on revealing spatial correlations by reconstructing masked patches within a single image, as seen in methods like SimMIM [27] and MAE [14]. These techniques are complementary. Methods such as SiameseIM [28] and MimCo [29] successfully integrate both CL and MIM to capture data features, leading to superior performance in both linear probing and fine-tuning evaluations.
Figure 1: Pretraining Workflow of the SeaMo Foundation Model. The SeaMo architecture integrates three primary components: encoders, temporal-multimodal fusion blocks (TM blocks), and decoders. Our approach incorporates a partially overlapping cropping strategy, ensuring that images from the same temporal instance are cropped identically across various modalities, while images from different instances exhibit partial overlaps. These processed images then serve as inputs to the network. Following the Masked Autoencoder paradigm, only visible tokens are processed by the encoder. The TM block effectively merges features from multiple seasons and modalities, culminating in modality-specific decoders that reconstruct the initially masked regions of the images.
@@ 2.2 Masked autoencoders
Masked autoencoders (MAE) [14] are a variant of denoising autoencoders that learn representations by reconstructing the original image from its masked inputs. This method is inspired by the self-masking pretraining mechanism found in the BERT[30] language model. It introduces random masks into the image inputs, requiring the network to reconstruct the original image from a partially masked version as a means of pretraining on image data. Typically, the model employs an encoder-decoder structure, and after pretraining, only the encoder is used as the backbone for downstream tasks. In masked image modeling, the encoder-decoder structure is usually asymmetric: the encoder is designed to learn high-level representations, while the decoder focuses on learning how to reconstruct the target. MAEs have proven effective across various vision benchmarks, prompting their adaptation in numerous studies across several data modalities. These include multimodal images [31], video data [32], medical data [33], remote sensing data [9] and meteorological data.[34].
Figure 2: Different crop strategies for seasonal data. The solid boxes indicate the images cropped and input into the network.(a) Images from different seasons are cropped to the same section. (b) Images from different seasons are cropped according to a specific proportion of the full image but ensure that some areas overlap. (c) Images from different seasons are cropped without any overlap.
@@ 3 Proposed Method
We aim to develop a foundation model, SeaMo, designed to learn robust temporal and spatial representations from multi-seasonal and multimodal remote sensing data. SeaMo consists of three principal components: encoders, temporal-multimodal fusion blocks, and decoders, as illustrated in Figure 1. Our research utilizes the SSL4EO-S12 dataset [35], which includes 12-channel multispectral optical imagery from Sentinel-2 and 2-channel SAR backscatter data from Sentinel-1.
3.1 Method overview of SeaMo
3.1.1 Tokenizer
We denote optical inputs and SAR inputs as $\bm{I}_{O},\bm{I}_{R}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ represents the number of images in a temporal sequence. The temporal image pairs are fed into the model season by season, denoted separately as $\bm{I}_{O}^{t},\bm{I}_{R}^{t}\in\mathbb{R}^{C\times H\times W},t=1,2,...,T$. Each type of data undergoes a modality-specific tokenizer into $L$ encodings $\bm{{E}}_{O}^{t},\bm{{E}}_{R}^{t}\in\mathbb{R}^{L\times D},t=1,2,...,T$ with fixed sinusoidal positional embeddings [36].
3.1.2 Mask operation

### v2

# SeaMo: A Season-Aware Multimodal Foundation Model for Remote Sensing

@@ Abstract
Remote Sensing (RS) data encapsulates rich multi-dimensional information essential for Earth observation. Its vast volume, diverse sources, and temporal continuity make it particularly well-suited for developing large Visual Foundation Models (VFMs). These models serve as powerful feature extractors, leveraging extensive RS data for pretraining and subsequent fine-tuning in various geoscientific applications. However, existing VFMs in the RS domain often concentrate on specific image characteristics, neglecting the full season-aware potential of RS data. To bridge this gap, we introduce SeaMo, a novel VFM that effectively integrates multimodal and multi-seasonal RS information. SeaMo leverages a masked image modeling framework to fully exploit the spatial, spectral, and seasonal dimensions of RS data. Specifically, we employ unaligned spatial region selection to capture spatial heterogeneity, incorporate multi-source inputs for enhanced multimodal integration, and introduce temporal-multimodal fusion blocks to assimilate seasonal variations effectively. By explicitly modeling the complex, season-dependent attributes of RS data, SeaMo enhances generalization, robustness, and adaptability across geoscientific tasks. Extensive experiments and ablation studies demonstrate its superior performance, underscoring its potential as a foundational model for Earth observation.
@@ 1 Introduction
Remote sensing is a technology that enables the measurement of soil or crop characteristics using platforms such as Unmanned Aerial Vehicles (UAVs), airplanes, or satellites. Data acquired through remote sensing technology are utilized to study social phenomena, land use dynamics, and changes in landscapes for scientific analysis and modeling purposes [1]. Currently, hundreds of remote sensing satellites continuously monitor the surface of the Earth, generating large-scale time series datasets. These datasets, characterized by their heterogeneity and incorporation of multiple sources, including spectral data, radar data, and meteorological data, provide a rich repository of physical geographic information, embodying the typical traits of big data [2]. The research on image data in remote sensing currently faces several challenges and focal points:
@@
(1) Integration of data from diverse sources for varied geoscientific applications: Different methods of capturing how electromagnetic radiation interacts with Earth’s surface materials provide unique data sets [3]. Optical data, by measuring radiation across many wavelengths, offers detailed spectral information that characterizes the material composition of objects. Synthetic Aperture Radar (SAR) data, by transmitting and receiving longer wavelength electromagnetic pulses, can assess the geometry, roughness, and electrical properties of objects. Given the significant heterogeneity between these data sources, simple transformations and combinations of data dimensions are insufficient to fully exploit this information [4]. Thus, a key area of focus is how to extract valuable geoscientific information from these diverse datasets that can meet human needs.
@@
(2) Effective modeling and utilization of spatio-temporal data with seasonal awareness: Satellite imagery inherently possesses a spatio-temporal structure as it consists of multiple observations of the same location over time. This rich temporal information enables the analysis of changes in physical geography, land use, and even offers insights into geological structures or socio-economic conditions [5]. However, the low information density resulting from low spatial resolution and long intervals between observations poses challenges for capturing temporal dynamics with simple techniques. Moreover, many existing methods overlook the critical impact of seasonal variations. Recognizing that seasonal changes play a fundamental role in shaping remote sensing observations, it becomes essential to develop models that are explicitly season-aware.
@@
VFMs refer to large-scale neural networks that are pretrained on vast datasets, typically using self-supervised learning methods. VFMs act as robust and powerful feature extractors and can be fine-tuned with minimal effort for various downstream tasks [6]. Given the inherent big data characteristics and the complex information contained within remote sensing data, the vast data landscape of remote sensing offers significant opportunities for the implementation of VFMs in Earth observation. Currently, several VFMs have been introduced in the field of remote sensing, demonstrating performance that surpasses traditional deep learning models on a broad range of geoscientific tasks. This advancement has led to significant breakthroughs in the mining and extraction of remote sensing information. However, these models still face several challenges and deficiencies:
@@
(1) Lack of geoscientific attributes: Some remote sensing VFMs are derived from computer vision technologies and often involve simplistic transformations of remote sensing data to adapt it to algorithmic requirements. This approach generally leads to insufficient exploration of the geoscientific properties inherent in remote sensing data, resulting in a gap in feature extraction capabilities. For instance, DINO-MC [7] employs basic data augmentation and uses Contrastive Learning (CL) for pretraining. Similarly, RSVA [8] straightforwardly adapts the Masked AutoEncoders (MAE) from the computer vision domain for pretraining in remote sensing.
@@
(2) Single-dimension data modeling: Although most remote sensing VFMs acknowledge the unique characteristics of remote sensing data, they typically focus on a single dimension, such as spatial or temporal attributes [9]. For instance, SeCo [10] and CaCo [11] exploit the temporal dynamics of RS images, employing CL to capture time-invariant features at specific geo-locations. RingMo [12] introduces Masked Image Modeling (MIM) to the RS domain, targeting small objects with a sparse masking strategy. SatMAE [13] employs a group combination masking strategy within the MAE framework [14], enriching its handling of multi-spectral or temporal data. Similarly, SpectralGPT [2] innovatively addresses the spectral sequence attributes of images, employing a 3D mask strategy focused on the reconstruction of spectral sequences. ScaleMAE [15] and SatMAE++ [16] adjust to the varying sizes and resolutions of RS images, designing multiple image-size reconstruction targets. CROMA [17] pretrains aligned radar and optical images using both CL and MAE methods.
@@
(3) Non-explicit modeling multi-dimensional data attributes: Some remote sensing VFMs consider multi-dimensional attributes, but their approach to modeling and pretraining is more akin to stacking and assembling rather than explicitly modeling from a remote sensing and geoscientific perspective [18, 19]. This often leads to a superficial integration of multidimensional data, which does not fully exploit the synergistic potential of combining spatial, temporal, and spectral characteristics inherent in remote sensing data. For instance, Skysense [18] integrates information from various RS data modalities, temporal information, and geographic location within the framework of CL. It employs CL for multimodal spatial feature extraction and utilizes positional encodings to embed temporal and geographic information, surpassing previous models. However, the integration of its multi-dimensional attributes is not fully realized. Specifically, the model optimizes the data learning from multiple separate aspects rather than in a holistic manner. The embedding of positional encodings serves merely as an insertion of prior knowledge rather than facilitating active knowledge learning and discovery.
@@
Motivated by these challenges, we argue that it is essential to develop a VFM that not only leverages the extensive, multi-source satellite data but also explicitly models both spatial and temporal properties under a unified training objective. To this end, we introduce SeaMo, a season-aware multimodal foundation model for remote sensing that effectively integrates spatial, multimodal, and multi-seasonal information. First, to enhance the spatial information density, we employ unaligned spatial region selection techniques. Importantly, these techniques are not merely used as data augmentation tools; they play a critical role in promoting the discovery of spatially diverse features and in challenging the model to learn more robust representations. Second, to harness the complementary strengths of multimodal satellite data during pretraining, we utilize a unified encoder that jointly accepts multisource data, with integration facilitated via self-attention mechanisms. Third, to improve time-invariant representations, we design a temporal fusion block that leverages cross-attention mechanisms. This block is more than a standard cross-attention module; it is purpose-built to fuse multiple spatial-temporal data streams, thereby capturing intricate relationships across both time and modality. Finally, the model is optimized using a unified masked modeling approach. Our contributions are outlined as follows:
@@
    We introduce SeaMo, a novel multimodal foundation model that leverages multi-season temporal and multimodal heterogeneous information embeddings. SeaMo integrates unaligned spatial region selection techniques to improve spatial diversity and model robustness effectively.
@@
    We design a temporal-multimodal fusion block, a cascade-style module that effectively integrates seasonal and multimodal information during pretraining. By employing a specialized form of cross-attention, this block goes beyond traditional mechanisms to capture intricate temporal-spatial interactions.
@@
    Our pretraining strategy adopts a multi-stage progression, evolving from unimodal learning to multimodal and ultimately seasonal-multimodal representation learning. This staged approach enables a more comprehensive understanding of the complex, multidimensional features characteristic of remote sensing data.
@@
    The proposed foundation model has been effectively transferred to a range of downstream geoscientific applications across diverse modalities, consistently demonstrating strong performance. Comprehensive ablation studies further validate its superior capabilities. By explicitly embedding seasonal awareness into multimodal feature extraction, SeaMo marks a substantial advancement in the development of foundation models for Earth observation.
@@ 2.1 Self-supervised visual representation learning
Self-supervised learning is a technique for learning generalizable visual representations through various pretext tasks without the need for annotations during pretraining [6]. Initially, this approach utilized straightforward tasks such as image colorization, image rotation prediction, or solving jigsaw puzzles [20, 21]. In recent years, CL and MIM have become predominant self-supervised techniques for pretraining large foundation models. CL is designed to capture feature invariance in a batch of images by differentiating between identical and modified images using data augmentation techniques, with implementations like SimCLR [22] and MoCo [23]. In contrast, MIM focuses on revealing spatial correlations by reconstructing masked patches within a single image, as seen in methods such as SimMIM [24] and MAE [14]. These techniques are complementary. Methods such as SiameseIM [25] successfully integrate CL and MIM to capture data features, leading to superior performance on both linear-probing and fine-tuning tasks.
Figure 1: Pretraining workflow of the SeaMo foundation model. The SeaMo architecture integrates three primary components: encoders, Temporal-Multimodal fusion blocks (TM blocks), and decoders. Our approach incorporates a partially overlapping spatial selecting strategy, ensuring that images from the same temporal instance are selected identically across various modalities, while images from different instances exhibit partial overlaps. These processed images then serve as inputs to the network. Following the masked autoencoder paradigm, only visible tokens are processed by the encoder. The TM block effectively merges features from multiple seasons and modalities, culminating in modality-specific decoders that reconstruct the initially masked regions of the images.
@@ 2.2 Masked autoencoders
MAEs [14] are a variant of denoising autoencoders that learn representations by reconstructing the original image from its masked inputs. It introduces random masks into the image inputs, requiring the network to reconstruct the original image from a partially masked version as a means of pretraining on image data. Typically, the model employs an encoder-decoder structure, and after pretraining, only the encoder is used as the backbone for downstream tasks. In MIM, the encoder-decoder structure is usually asymmetric: the encoder is designed to learn high-level representations, while the decoder focuses on learning how to reconstruct the target. MAEs have demonstrated their effectiveness on a wide range of vision benchmarks, prompting their adaptation in numerous studies spanning various data modalities. These include multimodal images [26], video data [27], medical data [28], remote sensing data [9] and meteorological data [29].
Figure 2: Different region selection strategies for temporal data. The solid boxes indicate the image regions that are selected and fed into the network. (a) Images from different seasons are selected from the same section. (b) Images from different seasons are selected based on a specific proportion of the full image, ensuring partial overlap. (c) Images from different seasons are selected with no overlap.
2.3 Remote sensing foundation models
As foundational models in remote sensing continue to flourish, new directions are emerging in remote sensing image processing. In this field, the training data for foundational models has expanded from single-sensor data to multi-sensor and even multi-modal data. For example, GFM [30] and ScaleMAE [15] pretrain on high-resolution RGB images, while Cross-Scale MAE [31] combines CL with MAE for training on high-resolution RGB images. In contrast, Models like SpectralGPT [2] develop self-supervised strategies tailored to the spectral characteristics of multispectral imagery. Additionally, some models enhance their training data sources by embedding extensive geographical and temporal information [32, 33]. For instance, Skysense [18] integrates location and temporal cues into a multimodal CL framework, whereas CaCo [11] performs inter-temporal CL to significantly raise the performance ceiling. In the development of Vision-Language Models (VLMs), the primary focus has been on aligning text with remote sensing images to enhance image understanding and facilitate interaction between remote sensing images and Large Language Models (LLMs). Models such as RemoteCLIP [34] and GeoCLIP [35] achieve few-shot capabilities, while some VLMs fine-tune both VLMs and LLMs on remote sensing-specific instruction datasets, enabling these models to interpret remote sensing images and engage in multi-turn visual dialogue tasks [36]. The model proposed in this work specifically targets the explicit fusion of data from multiple sensor sources while embedding temporal information.
@@ 3 Proposed Method
We propose SeaMo, a novel season-aware multimodal foundation model for remote sensing that learns robust spatial and temporal representations from multi-seasonal, multisource data. SeaMo comprises both a deep network architecture and an integrated training strategy. The network architecture includes a Multimodal Patch Embedding Module that transforms raw optical and SAR images into token embeddings, a Unified Multimodal Feature Encoder that jointly accepts multi-source data and integrates it via self-attention mechanisms, a Temporal Fusion Module for Multimodal Integration that leverages cross-attention to fuse information across seasons and modalities, and Modality-Specific Reconstruction Modules that recover the original images through masked image modeling. Complementing this architecture, our training strategy incorporates Geospatial Region Selection Strategies to enhance spatial representation learning via diverse cropping methods, as well as a Progressive Pretraining Strategy that transitions the model from single-time point multimodal learning to a comprehensive multi-time flow framework. Figure 1 illustrates the overall SeaMo framework. Our approach is pretrained on the SSL4EO-S12 dataset [37], which includes 12-channel multispectral optical imagery from Sentinel-2 and 2-channel SAR backscatter data from Sentinel-1.
3.1 SeaMo overview
3.1.1 Multimodal patch embedding module
We denote the optical inputs as $\bm{I}_{O}\in\mathbb{R}^{T\times C_{O}\times H\times W},$ and the SAR inputs as $\bm{I}_{R}\in\mathbb{R}^{T\times C_{R}\times H\times W}.$ Here, $T$ represents the number of temporal instances (or seasons), $C_{O}$ and $C_{R}$ denote the number of channels for optical and SAR data respectively, $H$ and $W$ are the height and width of each image. For each season $t=1,2,\ldots,T$, the corresponding images are given by $\bm{I}_{O}^{t}\in\mathbb{R}^{C_{O}\times H\times W}\quad\text{and}\quad\bm{I}_{R}^{t}\in\mathbb{R}^{C_{R}\times H\times W}.$ Each modality-specific image is processed by its dedicated embedding module, which converts the image into a set of $L$ patch tokens. Specifically, for each season $t$, the embedding modules yield token encodings $\bm{E}_{O}^{t},\ \bm{E}_{R}^{t}\in\mathbb{R}^{L\times D},$ where $L$ is the number of tokens and $D$ is the embedding dimension. Fixed sinusoidal positional embeddings [38] are added to these token encodings to preserve spatial order.
3.1.2 Selective masking of patch tokens